OpenStudy (anonymous):

The midpoint of PQ is (–6, –3). One endpoint of PQ is P(–1, –4). What are the coordinates of Q? A. Q(–11, –2) B. Q(–3.5, –3.5) C. Q(2, 11) D. Q(11, 2)

OpenStudy (zzr0ck3r):

So you have \(\dfrac{-1+x_2}{2}=-6\) and \(\dfrac{-4+y_2}{2}=3\). Solve for \(x_2\) and \(y_2\) and you answer is \((x_2, y_2)\)

OpenStudy (anonymous):

So (-1, 2) and (-2,2) ?

OpenStudy (zzr0ck3r):

\(\dfrac{-1+x_2}{2}=-6\implies -1+x_2 = -12\implies x_2=-11 \)

OpenStudy (zzr0ck3r):

Do something similar to get \(y_2\)

OpenStudy (anonymous):

ooh so y= -2

imqwerty (imqwerty):



imqwerty (imqwerty):

a=-11 nd b=-2

OpenStudy (anonymous):

Thank you both, I have a ton of questions that involve distance formula and I am not good at any of them

imqwerty (imqwerty):

those questions will become easy (: u jst need to have your basic concepts clear.

OpenStudy (zzr0ck3r):

The midpoint between \((a,b)\) ad \((c,d)\) is \((\dfrac{a+c}{2},\dfrac{b+d}{2})\)